To keep homemade ice cream frozen, it is placed in a bucket of ice-water and salt mixture. A mixture is commonly composed of 1.90 kg of salt (NaCl) and 7.25 kg of ice.

Problem: Calculate the mole fraction of NaCl in the mixture until the ice melts.


1
Expert's answer
2022-01-31T20:26:01-0500

Molar mass of NaCl"=58.44g\/mol"

Molar mass of water"=18.01528g\/mol"


Moles of NaCl"=\frac{1.90}{58.44}=0.032512"

Moles of Ice "=\frac{7.25}{18.01528}=0.402436"


Total fraction"=\frac{0.032512}{(0.032512+0.402436)}"

"=0.074749"

"=7.4749\%"